OpenStudy (anonymous):

Show that the polynomial \(x^4+x^2+1\) has no integer roots, but that it has a root modulo 3, and factorize it over \(Z_3\)

OpenStudy (anonymous):

By the rational roots theorem the only possible roots for x^4+x^2+1 = 0 are 1 or -1, and a manual check shows neither works. If you don't want to reference the rational roots theorem, you can rewrite the equation as x^2(x^1+1) = -1, which shows x is a divisor of 1. Clearly x = 1 and -1 are both roots mod 3, (+/-1)^4+(+/-1)^2+1 = 1+1+1 = 0 mod 3. So x^2-1 divides x^4+x^2+1 = (x^2-1)^2 mod 3.

OpenStudy (anonymous):

i get the second half but i'm still a little confused about the first half... in the book, it only shows \(x^4+x^2+1=(x^2+2)(x^2+2)=(x+1)(x+1)(x+2)(x+2)\) which does not make any sense -_-

OpenStudy (anonymous):

@love_jessika15 can you explain more?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

You got me confused now lol

OpenStudy (anonymous):

sorry... i'm so confused as well :(

OpenStudy (anonymous):

@satellite73 can you help?

OpenStudy (anonymous):

it's ok and thanks

OpenStudy (anonymous):

@jim_thompson5910 can you help me?

jimthompson5910 (jim_thompson5910):

(x^2+2)(x^2+2) = x^4 + 2x^2 + 2x^2 + 4 (x^2+2)(x^2+2) = x^4 + 4x^2 + 4 when you reduce the right side mod 3, you basically take the remainders if you divide each coefficient by 3 so 1 = 1 (mod 3), 4 = 1 (mod 3) because 4/3 = 1 remainder 1 which means (x^2+2)(x^2+2) = x^4 + 4x^2 + 4 = x^4 + x^2 + 1 (mod 3)

OpenStudy (anonymous):

do I do the same thing to get \((x+1)^2(x+2)^2\)?